# The $u$-invariant of function fields in one variable

### v1

@@ 6. Function fields in one variable over a henselian valued field
Assume that $v$ is a complete rank-$1$ valuation on $K$ such that $\mathsf{char}(Kv)\neq 2$ and let $r=\mathsf{rrk}(v)$. In this setting, in [21, Theorem 4.1] (independently from this work) the bound $\widehat{u}(K)\leqslant 2^{r}\widehat{u}(Kv)$ is shown, which strengthens [23, Cor. 6.2], where the same bound is given for the special case where $vK\simeq\mathbb{Z}^{r}$, while otherwise achieving the weaker bound $\widehat{u}(K)\leqslant 2^{r+1}\widehat{u}(Kv)$.
[…]
@@
[…]
@@
Let $k$ be a field with $\mathsf{char}(k)\neq 2$ and $\Gamma$ an ordered abelian group. Let $K$ denote the field of Hahn series with respect to $\Gamma$ and with coefficients in $k$. (The elements of $K$ are series $\sum_{\gamma\in S}a_{\gamma}t^{\gamma}$ with well-ordered support $S\subseteq\Gamma$.) Then $K$ carries a natural valuation $v$ with value group $\Gamma$ and residue field $k$, and $(K,v)$ is henselian, by [22, Lemma 2.40]. By 6.2, we obtain that $\widehat{u}(K)=[\Gamma:2\Gamma]\cdot\widehat{u}(k)$.
@@
Consider a valued field $(k,v)$ with $\mathsf{char}(kv)\neq 2$ and value group $vk$ contained in $\mathbb{R}$ as an ordered group. In particular $\mathsf{rk}(v)\leqslant 1$. Let $G$ denote the divisible closure of $vk$, which is still a subgroup of $\mathbb{R}$, and in fact a $\mathbb{Q}$-linear subspace of $\mathbb{R}$. We consider the quotient space $\mathbb{R}/G$. Let $n\in\mathbb{N}$ and choose $\gamma_{1},\dots,\gamma_{n}\in\mathbb{R}$ such that $\overline{\gamma}_{1},\dots,\overline{\gamma}_{n}$ are $\mathbb{Q}$-linearly independent in $\mathbb{R}/G$. Let $K=k(X_{1},\dots,X_{n})$. By [13, Theorem 2.2.1], $v$ extends to a valuation $w$ on $K$ such that $w(X_{i})=\gamma_{i}$ for $1\leqslant i\leqslant n$ and with $Kw=kv$ and $wK\simeq vk\times\mathbb{Z}^{n}$. In particular, we have $[wK:2wK]=2^{n}\cdot[vk:2vk]$.
@@
Let $k$ be a field in which $-1$ is not a square. Let us fix coordinate functions ${X,Y}$ on $\mathbb{P}_{k}^{2}$. Let $D=\pi^{-1}((X^{2}+1))$, the divisor of $\mathbb{P}_{k}^{2}$ induced by the fiber over $(X^{2}+1)$ of the projection $\pi:\mathbb{P}_{k}^{2}\rightarrow\mathbb{P}_{k}^{1}$, $(X,Y)\mapsto X$. We note that $D$ can be identified to $\mathbb{P}_{k(i)}^{1}$, where $i=\sqrt{-1}$ and $k(i)$ is the residue field of $(X^{2}+1)$ in $\mathbb{P}_{k}^{1}$. Its generic point $\eta_{D}$ corresponds to the ideal $(X^{2}+1)$ in $k[X,Y],$ so the local ring $\mathcal{O}_{\mathbb{P}_{k}^{2},\eta_{D}}$ is $k[X,Y]_{(X^{2}+1)}$, and it is a discrete valuation ring. Let $v_{1}$ be the discrete valuation it induces on $k(X,Y)$. The corresponding residue field is $k(i)(Y)$.
@@
We conclude by pointing out that in many cases when 6.2 is applied with a particular residue field $Kv$, on can obtain the precise value of $u(F)$ of every individual function field in one variable $F/K$.
Corollary 6.9.
Let $v$ be a henselian valuation on $K$ with $\mathsf{char}(Kv)\neq 2$. Let $n\in\mathbb{N}$ be such that $\widehat{u}(Kv)=2^{n}$. Then $\widehat{u}(K)=2^{n}\cdot[vK:2vK]$. Furthermore, if $u(\ell)=2^{n}$ for every finite field extension $\ell/Kv(\sqrt{-1})$, then for every function field in one variable $F/K$ we have $u(F)=2^{n+1}\cdot[vK:2vK]$.
Proof.
The first part is just a restatement of 6.2. To prove the second part, assume now that $u(\ell)=2^{n}$ for every finite field extension $\ell/Kv(\sqrt{-1})$. Consider a function field in one variable $F/K$. We choose $x\in F$ transcendental over $K$ and a discrete valuation $w$ on $F$ trivial on $K$ and with $w(x^{2}+1)>0$. Then $K(\sqrt{-1})\subseteq Fw$. We set $K^{\prime}=Fw$. Then $K^{\prime}/K$ is a finite field extension, and we obtain by 4.1 that $u(F)\geqslant 2u(K^{\prime})$. We denote by $v^{\prime}$ the unique extension of $v$ to $K^{\prime}$. Then $(K^{\prime},v^{\prime})$ is henselian and, by [13, Cor. 3.2.3], we have $[v^{\prime}K^{\prime}:vK]\cdot[K^{\prime}v^{\prime}:Kv]\leqslant[K^{\prime}:K]<\infty$. We conclude by 2.1 and 4.2, respectively, that
$[v^{\prime}K^{\prime}:2v^{\prime}K^{\prime}]=[vK:2vK]$ and $u(K^{\prime})=[v^{\prime}K^{\prime}:2v^{\prime}K^{\prime}]\cdot u(K^{\prime}v^{\prime}).$
As $K(\sqrt{-1})\subseteq Fw=K^{\prime}$, we have $Kv(\sqrt{-1})\subseteq K^{\prime}v^{\prime}$. Since $K^{\prime}v^{\prime}/Kv(\sqrt{-1})$ is a finite field extension, our assumption applies and yields that $u(K^{\prime}v^{\prime})=2^{n}$. Hence
$2^{n+1}\cdot[vK:2vK]=2u(K^{\prime})\leqslant u(F)\leqslant 2\widehat{u}(K)=2^{n+1}\cdot[vK:2vK]\,.$
Therefore $u(F)=2^{n+1}\cdot[vK:2vK]$. ∎
In the following cases, 6.9 provides directly the exact value of $u(F)$ for every function field in one variable $F/K$.
@@
Let $n\in\mathbb{N}$, $v$ a henselian valuation on $K$ with $[vK:2vK]=2^{n}$ and $F/K$ be a function field in one variable.

### v2

@@ 6. Function fields in one variable over a henselian valued field
Assume that $v$ is a complete rank-$1$ valuation on $K$ such that $\mathsf{char}(Kv)\neq 2$ and let $r=\mathsf{rrk}(vK)$. In this setting, in [26, Theorem 4.1] (independently from this work) the bound $\widehat{u}(K)\leqslant 2^{r}\widehat{u}(Kv)$ is shown, which strengthens [28, Cor. 6.2], where the same bound is given for the special case where $vK\simeq\mathbb{Z}^{r}$, while otherwise achieving the weaker bound $\widehat{u}(K)\leqslant 2^{r+1}\widehat{u}(Kv)$.
The equality $u(K)=[vK:2vK]\cdot u(Kv)$ obtained through 6.2 is an improvement to the bound $\widehat{u}(K)\leqslant 2^{r}\widehat{u}(Kv)$. Indeed, using 2.1 with $G=vK$ and $H=0$, we obtain that $vK:2vK]\leqslant 2^{r}$. Note also that this inequality can be strict (e.g. when $vK=\mathbb{R}$, we have $r=\infty$ and $[vK:2vK]=1$).
@@
Let $k$ be a field with $\mathsf{char}(k)\neq 2$ and let $K$ be the field of formal Puiseux series in one variable $t$ over $k$, that is, $K=\bigcup_{r\in\mathbb{N}}k(\!(t^{1/r})\!)$. The $t$-adic valuation on k((t))k(\!(t)\!) extends uniquely to a valuation $v$ on $K$, and this extension has residue field $Kv=k$ and value group $vK=\mathbb{Q}$. In particular, $\mathsf{rrk}(vK)=1$. Note that $v$ is henselian, but not complete. As $\mathbb{Q}$ is divisible, we have that $[vK:2vK]=1$. Hence we obtain by 6.2 that $\widehat{u}(K)=\widehat{u}(k)$. Denoting by $(K^{\prime},v^{\prime})$ the completion of $(K,v)$, we obtain similarly that $\widehat{u}(K^{\prime})=\widehat{u}(k)$.
@@
Let $k$ be a field with $\mathsf{char}(k)\neq 2$ and $\Gamma$ an ordered abelian group. Let $K$ denote the field of Hahn series with respect to $\Gamma$ and with coefficients in $k$. (The elements of $K$ are series $\sum_{\gamma\in S}a_{\gamma}t^{\gamma}$ with well-ordered support $S\subseteq\Gamma$.) Then $K$ carries a natural valuation $v$ with value group $\Gamma$ and residue field $k$, and $(K,v)$ is henselian, by [27, Lemma 2.40]. By 6.2, we obtain that $\widehat{u}(K)=[\Gamma:2\Gamma]\cdot\widehat{u}(k)$.
@@
Consider a valued field $(k,v)$ with $\mathsf{char}(kv)\neq 2$ and value group $vk$ contained in $\mathbb{R}$ as an ordered group. In particular $\mathsf{rk}(v)\leqslant 1$. Let $G$ denote the divisible closure of $vk$, which is still a subgroup of $\mathbb{R}$, and in fact a $\mathbb{Q}$-linear subspace of $\mathbb{R}$. We consider the quotient space $\mathbb{R}/G$. Let $n\in\mathbb{N}$ and choose $\gamma_{1},\dots,\gamma_{n}\in\mathbb{R}$ such that $\overline{\gamma}_{1},\dots,\overline{\gamma}_{n}$ are $\mathbb{Q}$-linearly independent in $\mathbb{R}/G$. Let $K=k(X_{1},\dots,X_{n})$. By [15, Theorem 2.2.1], $v$ extends to a valuation $w$ on $K$ such that $w(X_{i})=\gamma_{i}$ for $1\leqslant i\leqslant n$ and with $Kw=kv$ and $wK\simeq vk\times\mathbb{Z}^{n}$. In particular, we have $wK:2wK]=2^{n}\cdot[vk:2vk]$.
@@
Let $k$ be a field in which 1-1 is not a square. Let us fix coordinate functions X,Y{X,Y} on $\mathbb{P}_{k}^{2}$. Let $D=\pi^{-1}((X^{2}+1))$, the divisor of $\mathbb{P}_{k}^{2}$ induced by the fiber over $(X^{2}+1)$ of the projection $\pi:\mathbb{P}_{k}^{2}\rightarrow\mathbb{P}_{k}^{1}$, $(X,Y)\mapsto X$. We note that $D$ can be identified to $\mathbb{P}_{k(i)}^{1}$, where $i=\sqrt{-1}$ and $k(i)$ is the residue field of $(X^{2}+1)$ in $\mathbb{P}_{k}^{1}$. Its generic point $\eta_{D}$ corresponds to the ideal $(X^{2}+1)$ in $k[X,Y],$ so the local ring $\mathcal{O}_{\mathbb{P}_{k}^{2},\eta_{D}}$ is $k[X,Y]_{(X^{2}+1)}$, and it is a discrete valuation ring. Let $v_{1}$ be the discrete valuation it induces on $k(X,Y)$. The corresponding residue field is $k(i)(Y)$.
@@
In the following cases, 5.3 together with 6.2 provide directly the exact value of $u(F)$ for every function field in one variable $F/K$.
@@
Let $n\in\mathbb{N}$, $v$ a henselian valuation on $K$ with $vK:2vK]=2^{n}$ and let $F/K$ be a function field in one variable.